OpenStudy (anonymous):

Find the zeros of f(x) = -x^2 + 4x - 3 by using a table and a graph. Enter the missing numeral of the answer 1, ___. Use a number only. Find the zeros of the function f(x) = x^2 - 5x - 6 by factoring. What is the largest zero of the function? Enter a number only. Find the zeros of the function g(x) = x^2 - 8x by factoring. What is the smallest zero of the function? Enter a number only. Find the zeros of the function h(x) = x^2 +4x by factoring. What is the smallest zero of the function? Enter a number only. Find the zeros of the function f(x) = x^2 - 7x + 6 by factoring. What is the largest zero of the function? Enter a number only. Find the zeros of the function g(x) = x^2- 6x - 16 by factoring. What is the largest zero of the function? Enter a number only.

OpenStudy (anonymous):

For the first part you have to plug in 1 for x and calculate the corresponding y value -(1)^2+4(1)-3=y=0 For the second part the factor is (x-6)(x+1)=0 so x=6,-1 (6 is the larger of these values) For the third part the factor is (x-8)(x+0)=0 so x=8,0 (0 is the smaller of these values) For the fourth part the factor is (x+4)(x+0)=0 so x=-4,0 (-4 is the smaller of these values) For the fifth part the factor is (x-1)(x-6)=0 so x=1,6 (6 is the larger of these values) For the sixth part the factor is (x-8)(x+2)=0 so x=8,-2 (8 is the larger of these values) You can check my factoring my multiplying each out and seeing if you get the original equation

OpenStudy (anonymous):

thanks